OpenStudy (anonymous):

Perform the indicated operation. Write your final answer exactly as a fraction in lowest terms. Please show all of your work. 4/8-2/9

OpenStudy (anonymous):

4/8 = 1/2 Therefore: 1/2 - 2/9 = 9/18 - 4/18 = 5/18